The results of a chemostat study on the growth of
Answer #1

We will assume that the given cell growth is following Monod equation and as there is no endogenous material so the Kd=0, μ=μnet,

μ=μmaxCs/Ks+Cs

As the system is a chemostat system μg=D,

Therefore D=μmaxCs/Ks+Cs

1/D=Ks+CsmaxCs

1/D=(Ksmax)(1/Cs)+(1/μmax)

Now try to plot the graph of 1/D against 1/Cs,

1/D hr

1/Cs g-1L

11.905

18.519

10

12.658

6.250

7.246

5.051

5.376

4.132

4.425

1/μmax=y-intercept=2

μmax=0.500 hr-1

Ksmax=slope=11.8-4.6/17.2-4.6=0.5714 gL-1.hr

Ks=0.5714 gL-1.hrx0.5 hr-1=0.2857 gL-1

rxg Cx= (μmaxCs/Ks+Cs)Cx

rx=(0.500Cs/0.2857+Cs)Cx

(B) To find the rate equation for the ethanol formation, assume its a growth associated product

(1/Cx)(dP/dt)=Yp/xμgrp=dP/dt=Yp/xμgCx

rp=Yp/x(μmaxCs/Ks+Cs)Cs

Now from the slope of graph of Cp against Cx, obtain Yp/x

From the graph we can see Yp/x=deltaP/deltax=slope

Yp/x=8.2-2.5/2.2-0.6=3.5625 g P.(gX)-1

rp= 3.5625(0.500Cs/0.2857+Cs)Cx